OpenStudy (anonymous):

A farmer wants to set aside a rectangular plot of land to contain 660 square meters. If the width is 7 less than the length, find the dimensions of the plot.

OpenStudy (anonymous):

call the length x and so the width is x-7 to get the area of a rectangle it's length times width so; area = 660 = x(x-7) => 0 = x^2 -7x -660 Then use the quadratic formula and use the positive value :) ... this should work

OpenStudy (anonymous):

ok great! I was completing the square and it wasn't working out i'll try the quadratic formula instead

OpenStudy (anonymous):

complete the square should work as well (x-7/2)^2 -49/4-660 = 0 (x-7/2)^2 = 2689/4 (x-7/2) = +- 2sqrt(165) x = 7/2 + 2sqrt(165)

OpenStudy (anonymous):

should get the same answer using each method

OpenStudy (anonymous):

thank you so much. i'll run the numbers again

OpenStudy (anonymous):

got it! may i ask you one more question?

OpenStudy (anonymous):

yh sure

OpenStudy (anonymous):

....